OpenStudy (anonymous):

The position of a particle in motion in the plane at time t is X(t) = −7 t I + sin(−3 t) J. At time any t, determine the following: (a) the speed of the particle is: ??? (b) the unit tangent vector to X(t) is: ??? I + ??? J (Note that, a unit vector is a vector whose length is 1. Multiplying any non-zero vector V by 1/|V| produces a unit vector, and multiplying any unit vector by -1 shows that there are two unit vector which are multiples of any non-zero vector. The unit tangent vector to X(t) is defined as V(t)/|V(t)|. )

OpenStudy (anonymous):

\[x(t) = (-7t) i + \sin(-3t)j\] (a) The velocity of the particle is the derivative with respect to time of the position, so take the derivative of x(t): \[v(t) = x'(t) = -7i + (-3)\cos(-3t)j\] [perhaps by "speed" they mean the absolute value of the velocity? Lets calculate that]: \[\sqrt{(-7)^2 + (-3 * \cos(-3t) )^2}\] \[\sqrt{(-7)^2 + (9 * \cos^2(-3t) )}\] (b) The tangent unit vector of the position is the velocity vector divided by the absolute value of the velocity vector, which is a mess: \[v(t) / \left| v(t) \right|\] \[v(t) = (-7/\sqrt{49+9\cos^2(-3t)}) i - 3\cos^2(-3t)/\sqrt{49+9\cos^2(-3t)} j\] Are you sure they didn't specify a value of t for the tangent vector calculation?

OpenStudy (anonymous):

says the J part of the equation is wrong !